OpenStudy (anonymous):

Without a calculator, evaluate tan^2(2π/3)

OpenStudy (usarmy3947):

tan(-2π/-3) = tan(2π/3) = sin(2π/3) / cos(2π/3) = (√3 / 2) / (-1 / 2) = - √3 finally tan(-2π/-3) = - √3

OpenStudy (usarmy3947):

i hope it's right

OpenStudy (anonymous):

Thanks :)

OpenStudy (usarmy3947):

no problem

OpenStudy (anonymous):

\[\tan \frac{ 2\pi }{3 }=\tan \left( \pi-\frac{ \pi }{ 3 } \right)=-\tan \frac{ \pi }{3 }=-\frac{ \frac{ \sqrt{3} }{2 } }{ \frac{ 1 }{2 }}=-\sqrt{3}\] \[\tan ^{2}\left( \frac{ 2\pi }{3 } \right)=\left( -\sqrt{3} \right)^{2}=3\]